# An iterative method for the solution of Laplace-like equations in high and very high space dimensions

### v1

@@ 1 Introduction
In a sense, the following two sections form the core of the present work. They deal with the distribution of the values $\|T^{t}\eta\|^{2}$ on the unit sphere. Section 3 treats the general case. We first show that the expected value $E$ and the variance $V$ of the expression $\|T^{t}\eta\|^{2}$, considered as a random variable, can be expressed in terms of the singular values of the matrix $T$. Using this representation, we show that for random matrices $T$ with assigned expected value $E=1$ the expected value of the corresponding variances $V$ not only tends to zero as the dimension $m$ goes to infinity, but also that these variances cluster the more around their expected value the larger $m$ is. Sampling large numbers of these variances supports these theoretical findings. We also study a class of matrices $T$ that are associated with interaction graphs. These matrices correspond to the case that some or all coordinate differences are introduced as additional variables and formed the original motivation for this work. The expected values $E$ that are assigned to these matrices always take the value one and the corresponding variances $V$ can be expressed directly in terms of the vertex degrees and the dimensions. Numerical experiments for standard classes of random graphs show that the variances tend to zero with high probability as the number of the vertices increases. As mentioned, the knowledge of the expected value and the variance assigned to a given matrix $T$ is of central importance since the distribution of the values $\|T^{t}\eta\|^{2}$ is usually not much different from the corresponding normal distribution. This can be checked experimentally by sampling these values for a large number of uniformly distributed points $\eta$ on the unit sphere.
@@ 2 Solutions as traces of higher-dimensional functions
$\int f(x)\,\mathrm{d}x\,=\,n\nu_{n}\int_{0}^{\infty}\!f(re)r^{n-1}\,\mathrm{d}r.$ (14)
@@
If the norm $|\!|\!|U|\!|\!|_{s}$ of $U$ is finite, then also the norm $\|U\|_{s}$ and
$\|U\|_{s}\leq|\!|\!|U|\!|\!|_{s}$ (18)
holds. This follows from the Cauchy-Schwarz inequality. Both norms scale in the same way. By (14), they coincide for radially symmetric functions.
@@
$\phi(\eta)=n\nu_{n}\int_{S^{n-1}}\bigg{|}\frac{\widehat{F}(r\eta)}{r^{2}}\bigg{|}\,r^{n-1}\,\mathrm{d}r$ (22)
@@ 3 A moment-based analysis
for a large number of vectors $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{n-m}$ with standard normally distributed components. The values $\|y\|^{2}$ are then distributed according to the $\chi^{2}$-distribution with $n-m$ degrees of freedom. The direct calculation of $\|y\|^{2}$ can thus be replaced by the calculation of a single scalar quantity following this distribution. The amount of work then remains strictly proportional to the dimension $m$, no matter how large the difference $n-m$ of the dimensions is. Let $T$ be the matrix assigned to the graph associated with the $\mathrm{C}_{60}$-fullerene molecule, which consists of of the ninety edges of a truncated icosahedron and its sixty corners as vertices. The degree of these vertices is three and the assigned variance therefore $V=9/380$. The frequency distribution of the values $\|T^{t}\eta\|^{2}$ for a million randomly chosen points $\eta$ on the unit sphere and the corresponding Gauss function (23) are shown in Fig. 1.
@@ 4 Orthogonal projections and measure concentration
By Lemma 8, we can restrict ourselves to the matrix $P$, which extracts from a vector in $\mathbb{R}^{n}$ its first $m$ components. We split the vectors in $\mathbb{R}^{n}$ into a part $x\in\mathbb{R}^{m}$ and a remaining part $y\in\mathbb{R}^{n-m}$. The set whose volume has to be calculated according to the above remark, then consists of the points in the unit ball for which
@@ 5 The iterative procedure
be the to the interval $\alpha\leq\lambda\leq b$ transformed Chebyshev polynomial $T_{k}$ of the first kind of degree $k$. The norm (16) of the iteration error (2) then satisfies the estimate
$\|U_{k}-U\|_{s}\leq\frac{2r^{k}}{1+2r^{2k}}\,\|U\|_{s}+\|U-U_{\rho}\|_{s},\quad r=\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1},$ (6)
where $\widehat{U}_{\rho}$ takes the same values as $\widehat{U}$ on the set of all $\omega$ for which
$1-\rho\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)$ (7)
@@
$|P_{k}(\lambda)|\leq 1\text{\;for $0\leq\lambda\leq a$},\quad|P_{k}(\lambda)|\leq\frac{2r^{k}}{1+2r^{2k}}\text{\;for $a\leq\lambda\leq b$}$
@@
Depending on the size of $\kappa=\|T^{t}\|^{2}/(1-\rho)$, the norm of the error $U-U_{k}$ soon reaches the size of the norm of $U-U_{\rho}$. The idea behind the estimate (6) is that in high space dimensions the condition (7) is satisfied for nearly all $\omega$ and that the part $U-U_{\rho}$ of $U$ is therefore often negligible. The set on which the condition (7) is violated is a subset of the sector
$S(\delta)=\big{\{}\omega\,\big{|}\|T^{t}\omega\|^{2}\leq\delta\|\omega\|^{2}\big{\}}$ (8)
for $\delta=1-\rho$. Let the expression $\|T^{t}\omega\|^{2}$ possess the angular distribution
@@
Consider the solution $U$ of the equation (9) for a rotationally symmetric right-hand side $F$ with finite norm $\|F\|_{s}$, and let $\chi(t)=1$ for $t\leq\delta$ and $\chi(t)=0$ for $t>\delta$. Then
@@
on a sufficiently large ball around the origin hold without losing too much. This applies in particular to approximations by sums of Gauss functions.
@@
$\phi(\eta)\,=\,n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r,$
@@
If one rewrites the integral still in terms of the distribution $f$, (16) follows. ∎
@@
This is the point where our considerations from the previous two sections apply. We have seen that the values $\|T^{t}\omega\|^{2}$ are approximately normally distributed, with a variance $V$ and a standard deviation $\sigma=\sqrt{V}$ that tend to zero in almost all cases as the dimensions increase. This justifies replacing the actual distribution (9) by the corresponding normal distribution with the density
@@
The first twelve $M_{k}=M_{k}(\sigma)$ assigned to the orthogonal polynomials (26) for the standard deviations $\sigma=1/16$, $\sigma=1/32$, and $\sigma=1/64$ are compiled in Table 1. They give a good impression of the speed of convergence that can be expected and are fully in line with our predictions.
$\sigma=1/16$ $\sigma=1/32$ $\sigma=1/64$
3.89105e-03 9.75610e-04 2.44081e-04
3.05162e-05 1.90734e-06 1.19209e-07
3.61818e-07 5.60431e-09 8.73754e-11
5.76547e-09 2.19991e-11 8.54318e-14
1.15764e-10 1.08156e-13 1.04465e-16
2.81207e-12 6.39348e-16 1.53362e-19
8.03519e-14 4.41803e-18 2.62800e-22
2.64591e-15 3.49602e-20 5.14916e-25
9.88479e-17 3.11842e-22 1.13557e-27
4.13835e-18 3.09685e-24 2.78394e-30
1.92237e-19 3.38976e-26 7.51127e-33
9.82752e-21 4.05580e-28 2.21192e-35
@@
For the matrices $T$ from Sect. 3 assigned to undirected interaction graphs, the angular distribution of the values $\|T^{t}\omega\|^{2}$ behaves very similarly to the case of the rescaled orthogonal projections. In fact, it differs in almost all cases only very slightly from a to the expected value one rescaled beta distribution. Recall that beta distributions depend on two parameters $\alpha,\beta>0$ and possess the density
@@
For the matrices $T$ assigned to interaction graphs, these quantities satisfy the condition (29) with very few exceptions, and the resulting densities (28) agree almost perfectly with the actual densities. The values (22) assigned to these densities thus reflect the convergence behavior presumably at least as well as the values resulting from the given approximation by a simple Gauss function. In a sense, this reverses the argumentation and emphasizes the role of the model problem analyzed in Sect. 4. From this point of view, the beta distributions are at the beginning and the normal distributions only the limit case for higher dimensions.
@@
For $h=1$ and summation indices $k$ ranging from $-2$ to $50$, the relative error is, for example, less than $0.0007$ on almost the whole interval $1\leq r\leq 10^{18}$, that is, in the per mill range on an interval that spans eighteen orders of magnitude. Such an accuracy is surely exaggerated in the given context, but the example underscores the outstanding approximation properties of the functions (34). Figure 4 depicts the corresponding function $\phi$. These observations are underpinned by the analysis of the approximation properties of the corresponding infinite series in (Scholz-Yserentant, , Sect. 5). It is shown there that these series approximate $1/r$ on the positive real axis with a relative error
@@
as $h$ goes to zero. Moreover, the functions $\phi$ and their series counterparts almost perfectly fulfill the equioscillation condition as approximations of the constant $1$, so that not much room for improvement remains. In this respect, the case considered here significantly differs from the maximum case studied in Braess-Hackbusch , Braess-Hackbusch_2 , and Hackbusch_2 , at the price of often considerably larger sums. High relative accuracy over a large interval requires more effort than high absolute accuracy.
@@
The practical feasibility of the approach depends on the representation of the tensors involved and the access to the Fourier transforms of the functions represented by them. An attractive choice are constructions based on Gauss-Hermite functions. Since the Fourier transforms of such functions are of the same type, the iterations do not lead out of this class. Approximations based on Gauss-Hermite functions are standard in quantum chemistry.
A central task not discussed here is the recompression of the data between the iteration steps in order to keep the amount of work and storage under control, a problem common to all tensor-oriented iterative methods. If one fixes the accuracy in the single coordinate directions, the process reduces to an iteration on the space of the functions defined on a given high-dimensional cubic grid, functions that are stored in a compressed tensor format. There exist highly efficient, linear algebra-based techniques for recompressing such data. A problem in the given context may be that the in this framework naturally looking discrete $\ell_{2}$-norm of the tensors does not match the underlying norms of the continuous problem. Another open question is the overall complexity of the process. A difficulty with our approach is that the given operators do not split into sums of operators that act separately on a single variable or a small group of variables, a fact that complicates the application of techniques such as of those in Bachmayr-Dahmen or Dahmen-DeVore-Grasedyck-Sueli . For more information on tensor-oriented solution methods for partial differential equations, see the monographs Hackbusch_1 and Khoromskij of Hackbusch and of Khoromskij and Bachmayr’s comprehensive survey article Bachmayr .

### v3

@@ 1 Introduction
In a sense, the following two sections form the core of the present work. They deal with the distribution of the values $\|T^{t}\eta\|^{2}$ on the unit sphere. Section 3 treats the general case. We first show that the expected value $E$ and the variance $V$ of the expression $\|T^{t}\eta\|^{2}$, considered as a random variable, can be expressed in terms of the singular values of the matrix $T$. Using this representation, we show that for random matrices $T$ with assigned expected value $E=1$ the expected value of the corresponding variances $V$ not only tends to zero as the dimension $m$ goes to infinity, but also that these variances cluster the more around their expected value the larger $m$ is. Sampling large numbers of these variances supports these theoretical findings. We also study a class of matrices $T$ that are associated with interaction graphs. These matrices correspond to the case that some or all coordinate differences are introduced as additional variables and formed the original motivation for this work. The expected values $E$ that are assigned to these matrices always take the value one and the corresponding variances $V$ can be expressed directly in terms of the vertex degrees and the dimensions. Numerical experiments for standard classes of random graphs show that the variances tend to zero with high probability as the number of the vertices increases. As mentioned, the knowledge of the expected value and the variance assigned to a given matrix $T$ is of central importance since the distribution of the values $\|T^{t}\eta\|^{2}$ is usually not much different from the corresponding normal distribution. This can be checked experimentally by sampling these values for a large number of points $\eta$ that are uniformly distributed on the unit sphere.
@@ 2 Solutions as traces of higher-dimensional functions
$\int_{\mathbb{R}^{n}}f(x)\,\mathrm{d}x\,=\,n\nu_{n}\int_{0}^{\infty}\!f(re)r^{n-1}\,\mathrm{d}r.$ (14)
@@
If the norm $|\!|\!|U|\!|\!|_{s}$ of $U$ is finite, then also the norm $\|U\|_{s}$ and $\|U\|_{s}\leq|\!|\!|U|\!|\!|_{s}$ holds. This follows from the Cauchy-Schwarz inequality. Both norms scale in the same way. By (14), they coincide for radially symmetric functions.
@@
$\phi(\eta)=n\nu_{n}\int_{0}^{\infty}\bigg{|}\frac{\widehat{F}(r\eta)}{r^{2}}\bigg{|}\,r^{n-1}\,\mathrm{d}r$ (21)
@@ 3 A moment-based analysis
for a large number of vectors $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{n-m}$ with standard normally distributed components. The values $\|y\|^{2}$ are then distributed according to the $\chi^{2}$-distribution with $n-m$ degrees of freedom. The direct calculation of $\|y\|^{2}$ can thus be replaced by the calculation of a single scalar quantity following this distribution. The amount of work then remains strictly proportional to the dimension $m$, no matter how large the difference $n-m$ of the dimensions is. Let $T$ be the matrix assigned to the graph associated with the $\mathrm{C}_{60}$-fullerene molecule, which consists of the ninety edges of a truncated icosahedron and its sixty corners as vertices. The degree of these vertices is three and the assigned variance therefore $V=9/380$. The frequency distribution of the values $\|T^{t}\eta\|^{2}$ for a million randomly chosen points $\eta$ on the unit sphere and the corresponding Gauss function (23) are shown in Fig. 1.
@@ 4 Orthogonal projections and measure concentration
By Lemma 8, we can restrict ourselves to the matrix $P$ that extracts from a vector in $\mathbb{R}^{n}$ its first $m$ components. We split the vectors in $\mathbb{R}^{n}$ into a part $x\in\mathbb{R}^{m}$ and a remaining part $y\in\mathbb{R}^{n-m}$. The set whose volume has to be calculated according to the above remark, then consists of the points in the unit ball for which
@@ 5 The iterative procedure
be the to the interval $a\leq\lambda\leq b$ transformed Chebyshev polynomial $T_{k}$ of the first kind of degree $k$. The norm (16) of the iteration error (2) then satisfies the estimate
$\|U_{k}-U\|_{s}\leq\frac{2r^{k}}{1+r^{2k}}\,\|U\|_{s}+\|U-U_{\delta}\|_{s},\quad r=\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1},$ (6)
where $\widehat{U}_{\delta}$ takes the same values as $\widehat{U}$ on the set of all $\omega$ for which
$\delta\leq\alpha(\omega)(\|T^{t}\omega\|^{2}+\mu)$ (7)
@@
$|P_{k}(\lambda)|\leq 1\text{\;for $0\leq\lambda\leq a$},\quad|P_{k}(\lambda)|\leq\frac{2r^{k}}{1+r^{2k}}\text{\;for $a\leq\lambda\leq b$}$
@@
Depending on the size of $\kappa=\|T^{t}\|^{2}/\delta$, the norm of the error $U-U_{k}$ soon reaches the size of the norm of $U-U_{\delta}$. The idea behind the estimate (6) is that in high space dimensions the condition (7) is satisfied for nearly all $\omega$ and that the part $U-U_{\delta}$ of the solution $U$ is therefore often negligible. The set on which the condition (7) is violated is a subset of the sector
$S(\delta)=\big{\{}\omega\,\big{|}\|T^{t}\omega\|^{2}\leq\delta\|\omega\|^{2}\big{\}}.$ (8)
Let the expression $\|T^{t}\omega\|^{2}$ possess the angular distribution
@@
Consider the solution $U$ for a rotationally symmetric right-hand side $F$ with finite norm $\|F\|_{s}$, and let $\chi(t)=1$ for $t\leq\delta$ and $\chi(t)=0$ for $t>\delta$. Then
@@
on a sufficiently large spherical shell around its origin hold. To see why, note that
$|\widehat{U}(\omega)|\leq\frac{1}{\delta}\,\frac{|\widehat{F}(\omega)|}{\|\omega\|^{2}}$ (14)
holds for the frequency vectors $\omega$ outside the sector (8). Therefore, the part of the solution associated with the $\omega\notin S(\delta)$ outside a sufficiently large ball around the origin can be neglected. In the high dimensions considered, the same holds for the part of the solution associated with the $\omega\notin S(\delta)$ in a sufficiently small ball around the origin, provided that the Fourier transform of $F$ can be controlled there.
@@
$\phi(\eta)\,=\,n\nu_{n}\int_{0}^{\infty}|\widehat{U}(r\eta)|\,r^{s+n-1}\,\mathrm{d}r$
@@
If one rewrites the integral still in terms of the density $f$, (17) follows. ∎
@@
This is where our considerations from the previous two sections come into play. We have seen that the values $\|T^{t}\omega\|^{2}$ are approximately normally distributed, with a variance $V$ and a standard deviation $\sigma=\sqrt{V}$ that tend to zero in almost all cases as the dimensions increase. This justifies replacing the actual distribution (9) by the corresponding normal distribution with the density
@@
The first twelve $M_{k}=M_{k}(\sigma)$ assigned to the orthogonal polynomials (27) for the standard deviations $\sigma=1/16$, $\sigma=1/32$, $\sigma=1/64$, and $\sigma=1/128$ are compiled in Table 1. They give a good impression of the speed of convergence that can be expected and are fully in line with our predictions.
$\sigma=1/16$ $\sigma=1/32$ $\sigma=1/64$ $\sigma=1/128$
3.891051e-03 9.756098e-04 2.440810e-04 6.103143e-05
3.051618e-05 1.907343e-06 1.192093e-07 7.450581e-09
3.618180e-07 5.604306e-09 8.737544e-11 1.364492e-12
5.765467e-09 2.199911e-11 8.543183e-14 3.332296e-16
1.157645e-10 1.081565e-13 1.044654e-16 1.017371e-19
2.812071e-12 6.393485e-16 1.533625e-19 3.727770e-23
8.035187e-14 4.418030e-18 2.627999e-22 1.593745e-26
2.645905e-15 3.496016e-20 5.149158e-25 7.788140e-30
9.884788e-17 3.118421e-22 1.135567e-27 4.282074e-33
4.138350e-18 3.096852e-24 2.783941e-30 2.616285e-36
1.922373e-19 3.389755e-26 7.511268e-33 1.758579e-39
9.827520e-21 4.055803e-28 2.211916e-35 1.289675e-42
@@
For the matrices $T$ from Sect. 3 assigned to undirected interaction graphs, the angular distribution of the values $\|T^{t}\omega\|^{2}$ is often very similar to that in the case of rescaled orthogonal projections. In fact, it differs in very many cases only very slightly from a to the expected value one rescaled beta distribution. Recall that beta distributions depend on two parameters $\alpha,\beta>0$ and possess the density
@@
For the matrices $T$ assigned to large classes of graphs, these quantities satisfy the condition (30) in general or with very few exceptions and the resulting densities (29) agree almost perfectly with the actual densities. The values (23) assigned to these densities thus reflect the convergence behavior presumably at least as well as the values resulting from the given approximation by a simple Gauss function. In a sense, this reverses the argumentation and emphasizes the role of the model problem analyzed in Sect. 4. From this point of view, the beta distributions are at the beginning and the normal distributions only the limit case for higher dimensions.
@@
For $h=1$ and summation indices $k$ ranging from $-2$ to $50$, the relative error is, for example, less than $0.0007$ on almost the whole interval $1\leq r\leq 10^{18}$, that is, in the per mill range on an interval that spans eighteen orders of magnitude. Such an accuracy is surely exaggerated in the given context, but the example underscores the excellent approximation properties of the functions (35). Figure 4 depicts the corresponding function $\phi$. These observations are underpinned by the analysis of the approximation properties of the corresponding infinite series in (Scholz-Yserentant, , Sect. 5). It is shown there that these series approximate $1/r$ on the positive real axis with a relative error
@@
as $h$ goes to zero. High relative accuracy on large intervals $1\leq r\leq R$ is a considerably stronger requirement than high absolute accuracy. The approximation of $1/r$ with minimum absolute error has been studied in Braess-Hackbusch , Braess-Hackbusch_2 , and Hackbusch_2 . The technique in Hackbusch_2 can be used to compute the approximations with the least relative error. On the given interval, one gains a factor of just over five with a sum of $50$ exponential functions. Hackbusch has compiled the best approximations for a large number of intervals and sums of exponential functions of very different length. The data can be downloaded from the website of the Max Planck Institute in Leipzig Hackbusch_3 .
@@
The practical feasibility of the approach depends on the representation of the tensors involved and the access to the Fourier transforms of the functions represented by them. A central task not discussed here is the recompression of the data between the iteration steps in order to keep the amount of work and storage under control, a problem common to all tensor-oriented iterative methods. If one fixes the accuracy in the single coordinate directions, the process reduces to an iteration on the space of the functions defined on a given high-dimensional cubic grid, functions that are stored in a compressed tensor format. There exist highly efficient, linear algebra-based techniques for recompressing such data. A problem in the given context may be that the in this framework naturally looking discrete $\ell_{2}$-norm of the tensors does not match the underlying norms of the continuous problem. Another open question is the overall complexity of the process. A difficulty with our approach is that the given operators do not split into sums of operators that act separately on a single variable or a small group of variables, a fact that complicates the application of techniques such as of those in Bachmayr-Dahmen or Dahmen-DeVore-Grasedyck-Sueli . For more information on tensor-oriented solution methods for partial differential equations, see the monographs Hackbusch_1 and Khoromskij of Hackbusch and of Khoromskij and Bachmayr’s comprehensive survey article Bachmayr .